OpenStudy (anonymous):

Find an equation of the tangent line to the graph of the function f(x) = 2sinx at the point where x = pi/3.

OpenStudy (anonymous):

ok in order to find tangent line we have to take the derivative and so we will find the slope of the tangent line what is the derivative of f(x) = 2sin(x) ?

OpenStudy (anonymous):

2cosx

OpenStudy (anonymous):

that's good.. f'(x) = 2cos(x) now since we want the tangent to the function at the point x=pi/3 in order to find the slope of the tangent we have to plug x=pi/3 into the derivative f'(pi/3) = ?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

yes so we have a slope for the tangent line m= 1.

OpenStudy (anonymous):

the last thing that we will need is a point on the tangent line so we could find its equation since the tangent line touches the graph at x=pi/3 we can get a y value from the original function f(x). so we will have a point (pi/3,f(pi/3))

OpenStudy (anonymous):

so f(pi/3) = 2sin(pi/3) = sqrt(3) now we have a point (pi/3,sqrt(3)) and a slope m =1 using the formula for line equation y=y1 + m(x-x1) can you find the tangent line ?

OpenStudy (anonymous):

ill give the answer since i have to go but try it by yourself as well : y = sqrt(3) + x - pi/3 y= x +sqrt(3)-pi/3

OpenStudy (anonymous):

thanks! i got it :)